OpenStudy (anonymous):

A video game has been developed whereby the player must earn a certain number of experience points before progressing to the next level. On level 1, a player must earn 2 experience points. On level 2, a player must earn 4 experience points. On level 3, a player must earn 7 experience points. On level 4, a player must earn 11 experience points. Which recursive equation represents the pattern? an = an – 1 + n an = an – 1 – n an = an – 1 + 2n an = an – 1 – 2n I just need the answer please.

OpenStudy (anonymous):

i already answered this for you this site is for learning, not just answers

OpenStudy (anonymous):

Is it A?

OpenStudy (anonymous):

It's 2:45am I can't learn at this timeeeeee

OpenStudy (primeralph):

Ahhh I can answer this.

OpenStudy (anonymous):

answer it ^^

OpenStudy (primeralph):

Sorry, I'd like to help, but I'm not allowed to give you the answer.

OpenStudy (anonymous):

Then just take it step by step without any interruptions?

OpenStudy (anonymous):

I already did this problem for you last time, i'm confused what the confusion is o.0

OpenStudy (anonymous):

Your process confused me

OpenStudy (e.cociuba):

The answer is A. You can see by taking the difference between the experience required between two levels is n. For example exp needed for level 4 - exp needed for level 3 = 11 - 7 = 4.

OpenStudy (anonymous):

THANK YOU.

OpenStudy (e.cociuba):

Do you understand?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Well kinda, but not enough to do it for a different problem...........

OpenStudy (goformit100):

Ya I understood it.

OpenStudy (e.cociuba):

Successive differences are 2, 3, 4. For example, when n = 2, a(2) = a(1) + 2. So that's why ur answer would be A.

OpenStudy (anonymous):

A video game has been developed whereby the player must earn a certain number of experience points before progressing to the next level. On level 1, a player must earn 1 experience point. On level 2, a player must earn 4 experience points. On level 3, a player must earn 8 experience points. On level 4, a player must earn 13 experience points. Which recursive equation represents the pattern? an = an – 1 + n an = an – 1 + (n + 1) an = an – 1 + 2n an = an – 1 + (2n + 1) If it were like this how would you do it?

OpenStudy (anonymous):

I answered with C and got it wrong.

OpenStudy (anonymous):

Successive differences are 3, 4, and 5 \[a_{n-1} \]is the amount of experience you need to earn from the last level you were on so if you're on level 2\[a_2 = 4,\] and \[a_{2-1} = 1\]